I'm not sure how to do part B, please help, and show the work if you can!
6-3 A cross between an Hfr bacterial strain with genotype val met tyr is crossed with an F strain with genotype val met ty
Answer #1

Recombination frequency = number of recombinants / total number of progenies × 100

Number of recombinants = always in the minimum number

Total number of progenies = 552 + 63 + 15 = 630

Total number of recombinants = 63 + 15 = 78

Recombination frequency = 78 / 630 × 100 = 12.38%

Please rate high.